Home‎ > ‎Crew 164‎ > ‎

0223-DailySummaryReport

DAILY SUMMARY REPORT
SOL: 3
Person filling out Report: Mohammad IRANMANESH, XO
Summary Title: Science and pancakes
Mission Status: Main systems are nominal. Except hot water.
Sol Activity Summary: EVA #2 - Preparation work on Personal Asistant for Marsonauts - NorCal connection test - Work on monitoring of martian atmosphere - CRV repair after EVA - Change of the EVA planning for the next few days.
Look Ahead Plan: tomorrow we will get the supplies from the resupply ship and continue with EVA 3. 
Anomalies in work: Hot water is not working but might be fixed with the planned filter change.
Weather: Sunny but windy.
Crew Physical Status: Everyone feeling great today. No health issue anymore.
EVA: EVA to explore around the HAB + test of CRV + test of the EMUI smart glasses.
Reports to be filed: 
- Commander report
- Engineering report
- Journalist report 
- EVA #3 request
- EVA #2 report
Support Requested: